A light bulb is rated at 30 W when operated at 120 V. How much charge enters (and leaves) the light bulb in 1.0 min?

a. 17 C

b. 15 C

c. 14 C

d. 13 C

e. 60 C
Answer #1

here

P = V * I

I = P / V

I = 30 / 120

I = 0.25 amp

and then

a = C / s

0.25 C/s * 60 s/min = 15 C/min

therefore there is 15 C charge in 1 minute

the correct answer is b) 15C